∫ (x^2+y^2)/(x-y) dx =
1
Expert's answer
2020-04-07T13:00:04-0400

"\int\frac{x^2+y^2}{x-y}dx=\int(\frac{2y^2}{x-y}+x+y)dx=2y^2\log(x-y)+\frac{x^2}{2}+xy+const"